### v1

@@ Abstract
The detection of outliers is of critical importance in the assurance of data quality. Outliers may exist in observed data or in data derived from these observed data, such as estimates and forecasts. An outlier may indicate a problem with its data generation process or may simply be a true statement about the world. Without making any distributional assumptions, this paper proposes the use of loss functions to detect these outliers in panel data. An unsigned loss function is derived axiomatically. A signed loss function is developed to account for positive and negative outliers separately. In the case of nominal time an exact parametrization of the loss function is obtained. A time-invariant loss function permits the comparison of data at multiple times on the same basis. Several examples are provided, including an example in which the outliers are classified by another variable.
1 Introduction
@@
In general, an outlier is an observation which departs from the norm (however defined) in a set of observations. Outliers can indicate a problem with the data generation process (i.e., anomalies) or may be true, but unusual, statements about reality. [^2] This is similar to Hoaglin’s (1983, p. 39-40) use of “outside cutoffs” to identify “outside values.” Our purpose in detecting outliers is to determine whether there are problems with the data generation process. In terms of Barnett and Lewis (1994, p. 37), we are testing for discordancy. This paper specializes the problem of detecting outliers to panel data, such as estimates and forecasts. Panel data are cross-sectional time series, such as a time series of population estimates for a set of areas. [^3] The bidimensionality of data searched for outliers is not unique: DuMouchel (1999), Albert (1997), and Rudas, Clogg and Lindsay (1994) search for outliers in contingency tables. The contingency table approach differs in that time need not be a dimension and that parametric assumptions are made. Time may be either chronological or nominal. Nominal time indexes different sets of estimates or forecasts for the same cross-sectional units and chronological time. Time is nominal in this context because the different estimates sets have no natural ordering. Comparing cross-sectional estimates to their true values is an instance of nominal time.
Section 2 develops loss functions to compare two sets of values. No distributional assumptions are made, as the natures of the data generation processes are assumed unknown and nonidentical. [^4] This obviates the use of parametric techniques, in which observations are tested for departure from a predetermined, hypothesized distribution. Subsection 2.1 specifies the assumptions and develops a simple loss function that satisfies these assumptions and can be made to satisfy the property of the next sentence, when the base and future dates are exactly one period apart. This Subsection further quantifies the property that the loss associated with a given relative difference increases in the initial value. Subsection 2.2 generalizes the loss function to period lengths other than one, so that comparisons can be made across time. Subsection 2.3 develops the signed loss function, which incorporates the direction of the difference. Subsection 2.4 specializes the loss function to compare two sets of estimates for the same cross-sectional units and chronological time. Section 3 discusses some applications, including general usage of loss functions, parametrizing loss functions from preexisting outlier criteria, using loss functions with GIS and parametrizing a variant of the loss functions of Section 2 when the outliers are classified by another variable. These examples are based on actual Census Bureau applications. Section 4 concludes this paper.
2 The Loss Function
@@
This section describes the assumptions used to generate the loss function $L(F;B)$ and its variants, where $F$ is the future value and $B$ is the base period value. The loss function is the penalty associated with the difference between $F$ and $B$. Roughly speaking, the greater the difference between $F$ and $B$, the greater the loss. Initially, $F$ is assumed to be one period after $B$. After the necessary assumptions are made, on of the simplest forms of $L$ is specified. Restrictions on the values of the parameters of $L$ which make it increase in $B$ for a given relative difference are then specified. Subsection 2.1 axiomatically develops the simplest unsigned loss function L which satisfies these properties. Subsection 2.2 generalizes L to situations in which F and B may not be exactly one period apart. Subsection 2.3 introduces the signed loss function for cases in which the sign of the difference is an additional important criterion. Subsection 2.4 parametrizes L for comparing two sets of estimates of the same parameters. Throughout this paper, $B$ and $F$ are assumed positive. Zeroes, which frequently arise in the contexts in which the authors have worked, are either recoded to small values or omitted from the analysis.
2.1 The Unsigned Loss Function
@@
The unsigned loss function $L$ is constructed by specifying three assumptions. [^5] The exposition in this Subsection through Property 1 (except for the use of Lie symmetry) follows Coleman (2025). The first assumption is that $L$ is symmetric in the differences:
Assumption 1: $L(B+\epsilon,B)=L(B-\epsilon,B)$ for all $B.\epsilon\geq 0$.
@@
This assumption is not as innocuous as it looks. It is quite possible that, at least for some range of $B$, that positive and negative differences have differential impacts. However, the resulting asymmetry complicates the definition of $L$. Subsection 2.2 relaxes this assumption by developing the signed loss function, which allows the possibility of incorporating the direction of the difference $\epsilon$. The symmetry of $L$ allows us to use the equivalent notation $\ell(\epsilon,B)=L(F;B)$ where $\epsilon=|F-B|$.
@@
Assumption 2: $\partial\ell/\partial\epsilon>0$ for all $\epsilon\geq 0$.
@@
Finally, we want $L$, or, equivalently, $\ell$ , to decrease in $B$. This means that for a given value of $\epsilon$, the loss associated with it decreases with its associated initial value. This has two justifications. First, for example, a difference of 500 when the initial value is 1,000 is a whopping 50%, a highly significant difference. However, the same difference, when the initial value is 1,000,000 is akin to a roundoff error. Second, when performing estimates or taking samples, the coefficient of variation $\sigma/\mu$, where $\sigma$ is the standard deviation and $\mu$ is the expected value, decreases in $B$. One author’s (Coleman) experience is that all areas tend to have about the same roundoff errors. Again, these are proportionately greater in small areas. We state this formally as:
Assumption 3: $\partial\ell/\partial B<0$ , or, equivalently, $\partial L/\partial B<0$ for all $B>0$.
One of simplest functions which satisfies Assumptions 1–3 and admits Property 1 below is the Cobb-Douglas function
$\epsilon^{p}B^{q}$ (1a)
@@
$|F-B|^{p}B^{q}.$ (1b)
An observed pair $(F;B)$ is an outlier whenever $L(F;B)>C$, where $C$ is a predetermined critical value. [^6] CC can also be determined from the data by taking a predetermined quantile or a multiple of the interquartile range of $L$ (Tukey 1977). We will also refer to outliers as being critical. We will also refer to the equation $L(F;B)=C$ as the equation of criticality. The choice of $p,q$ and $C$ is an empirical matter. Only a practitioner’s experience with data can determine when data are suspect and incorporate these suspicions into parameters. One thing to note is that the loss function is ordinal: raising $L$ or $\ell$ and $C$ to any positive power $m$ leaves the rankings of losses unchanged. It is only the rankings of losses that are important. Formally, the transformation $p\to p_{m}=mp$, $q\to q_{m}=mq$ and $C\to C_{m}=mC$ creates a new equation of the same form. The same independent variables create the same criticality (or lack thereof). This is called a Lie symmetry or Lie invariance. The upshot is that one parameter can be fixed and the other two varied to generate new loss functions and critical values. In subsequent development, $p$ will be set equal to 1. Thus,
$\ell(F,B)=\epsilon B^{q}$ (2a)
@@
Property 1: The loss function defined by equations (2a) and (2b) increases in $B$ for any given absolute relative difference. This is assured whenever q>1q>-1.
The reader may notice that q = 0 turns the first terms of equations (2a) and (2b) into the absolute values of the differences. Thus, values of $q$ between 0 and 1–1 represent various tradeoffs of absolute differences and absolute relative differences. Consider the product of the $r$th power of the absolute difference and the $s$th power of the absolute relative difference, where $r,s>0$,
$|F-B|^{r}\left(\frac{|F-B|}{B}\right)^{s}.$ (4)
By the Lie symmetry invoked above, loss function (4) is isomorphic to the loss function $|F-B|B^{-s/r+s}$. Thus, any value of q corresponds to an infinite number of pairs $(r,s)$ where q=s/(r+s)q=–s/(r+s). Geometrically, the same loss function is generated for all $(r,s)$ lying on the line r=(1+q)sr=–(1+q)s.
Subsection 3.4 provides an example of a variant of this loss function approach when the outliers are classified by another variable. Changes in the variable of interest are declared outliers contingent on another variable. Assumption 3 is violated, but, in this example, Property 1 holds.
2.2 The Time-Invariant Loss Function
@@
$L(F_{it};B_{i},t)=\left(\frac{|F_{it}-B_{i}|}{B_{i}}\right)^{\frac{1}{t}}B_{i}^{q+1}=|F_{it}-B_{i}|^{\frac{1}{t}}B_{i}^{q+1+\frac{1}{t}}$ (5)
Given this paper’s framework, equation (5) should be used to make comparisons across time, as it puts the geometric average absolute relative difference on the same basis for all $t$.
Now, we can use Lie symmetry to make the exponent of $|F_{it}-B_{i}|$ unity by multiplying both exponents by $t$. We can also verify that the exponent of $B_{i}$ lies between 0 and –1:
$L(F_{it};B_{i},t)=|F_{it}-B_{i}|B_{i}^{tq+t-1}.$ (6)
@@
2.3 The Signed Loss Function
@@
The time-invariant signed loss function (after accounting for the Lie symmetry) is:
$S(F_{it};B_{i},t)=(F_{it}-B_{i})B_{i}^{tq+t-1}.$ (8)
2.4 Comparing Two Sets of Data: A Specialization of the Loss Function
Suppose that the sets $\mathbf{B}=\{B_{i}\}$ and $\mathbf{F}=\{F_{i}\}$ represent two versions of estimates of the true values $\mathbf{A}=\{A_{i}\}$. [^7] The derivations in this Subsection were done with assistance from ChatGPT 5. Suppose that both the $B_{i}$ and $F_{i}$ are unbiased estimators of the $A_{i}$ and that their variances are proportionate to the $A_{i}$ (i.e., $\operatorname{Var}(B_{i})=\operatorname{Var}(F_{i})=\sigma^{2}A_{i}$.)888 This can be justified by assuming that the $B_{i}$ and $F_{i}$ are both sums of $A_{i}$ independent (or, more weakly, jointly uncorrelated) estimators with expected value 1 and variance $\sigma^{2}$. Suppressing subscripts, consider the expected value of the squared difference of $B$ and $F$, $\operatorname{\mathbf{E}}(F–B)^{2}$:
$\operatorname{\mathbf{E}}(F-B)^{2}=\operatorname{Var}(F)+\operatorname{Var}(B)-2\operatorname{Cov}(F,B)+(\operatorname{\mathbf{E}}F-\operatorname{\mathbf{E}}B)^{2}$ (9)
@@ 2.4 Comparing Two Sets of Data: A Specialization of the Loss Function
$\operatorname{\mathbf{E}}(F-B)^{2}=\operatorname{Var}(F)+\operatorname{Var}(B)-2\operatorname{Cov}(F,B)$ (10)
@@
$\operatorname{\mathbf{E}}(F-B)^{2}=2\sigma^{2}A-2\operatorname{Cov}(F,B).$ (11)
@@
$\operatorname{Cov}(F,B)=\rho\operatorname{Var}(F)\operatorname{Var}(B)=\rho\sigma^{2}A.$ (12)
Substitutiong equation (12) into equation (11) yields
$\operatorname{\mathbf{E}}(F-B)^{2}=2(1-\rho)\sigma^{2}A\propto A.$ (13)
@@
$(F-B)^{2}\propto A=\operatorname{\mathbf{E}}B.$ (14)
@@
$|F-B|B^{-\frac{1}{2}}\propto 1.$ (16)
This means that, in this situation, we can use the loss functions (2a) and (2b) with q=1/2q=–1/2. Since the null distributions of B and F are assumed unknown, it is impossible to do any significance testing. Moreover, since we are dealing with the entire population, sampling theory breaks down. Even though estimates of $\rho$ and $\sigma$ can be computed from the data, it is unclear how to use normal (or any other distributional) theory for testing. The simplest thing to do in this situation is to rank the losses in descending order and empirically decide what the one critical value is.
Again, the signed loss function (7) can be used with q=1/2q=–1/2.
3 Applications
This section illustrates the use of loss functions by first outlining a general procedure for using loss functions in Subsection 3.1. Next, three different examples of loss functions are shown. In the first example, in Subsection 3.2, preexisting outlier criteria in terms of critical ratios by size class are transformed into a loss function. The second example, in Subsection 3.3, uses real-world data and GIS to compare the 2010 and 2020 Census couonty counts using the q=1/2q=–1/2 loss function of Subsection 2.4 as well as absolute and absolute relative differences to evaluate differences between these two sets of estimates. The third example, in Subsection 3.4, uses preexisting outlier criteria for a variable, which are classified by another variable, to generate a different type of loss function.
3.1 General Procedure for Using Loss Functions
Loss function evaluations usually begin by recoding zeroes to a small positive value (the exact value determined by the range and smallest value of the data) and setting q=.5q=-.5. If time is chronological, the analyst then has to examine the data and the rankings of their associated losses. If, in their opinion, too many observations with small changes occurring to small base values are ranked highly, then $q$ should be increased. If, on the other hand, too many observations with small changes to large base values are ranked highly, then $q$ should be decreased. This process continues until the analyst is satisfied with the loss rankings. The authors have found that initially changing $q$ by .1 is satisfactory. If time is nominal, the authors’ experience is that two comparisons should be done using both variables for $B$.

### v2

@@ Abstract
The detection of outliers is of critical importance in the assurance of data quality. Outliers may exist in observed data or in data derived from these observed data, such as estimates and forecasts. An outlier may indicate a problem with its data generation process or may simply be a true, but unusual, statement about the world. Without making any distributional assumptions, we proposes the use of loss functions to detect these outliers in panel data.
Part I covers nonnegative data. We axiomatically derive an unsigned loss function. We then develop a signed loss function ito account for positive and negative outliers separately. In the case of nominal time we obtain an exact parametrization of the loss function. A time-invariant loss function permits the comparison of data at multiple times on the same basis. We provide several examples, including an example in which the outliers are classified by another variable.
Part II covers data of mixed sign. Similar to Part I, we axiomatically develop unsigned and signed loss functions. We search for optimal values of the loss function parameter using graphs.
Introduction
@@
In general, an outlier is an observation which departs from the norm (however defined) in a set of observations. Outliers can indicate a problem with the data generation process (i.e., anomalies) or may be true, but unusual, statements about reality. [^2] This is similar to Hoaglin’s (1983, p. 39-40) use of “outside cutoffs” to identify “outside values.” Our purpose in detecting outliers is to determine whether there are problems with the data generation process(es). In terms of Barnett and Lewis (1994, p. 37), we are testing for discordancy. This paper specializes the problem of detecting outliers to panel data, such as estimates and forecasts. Panel data are cross-sectional time series, such as a time series of population or migration estimates for a set of areas. [^3] The bidimensionality of data searched for outliers is not unique: DuMouchel (1999), Albert (1997), and Rudas, Clogg and Lindsay (1994) search for outliers in contingency tables. The contingency table approach differs in that time need not be a dimension and that parametric assumptions are made. We make no distributional assumptions, as we assume the data generation processes to be unknown and nonidentical. [^4] This obviates the use of parametric techniques, in which observations are tested for departure from a predetermined, hypothesized distribution. Time may be either chronological or nominal. Nominal time indexes different sets of estimates or forecasts for the same cross-sectional units and chronological time. Time is nominal in this context because the different estimates sets have no natural ordering. Comparing cross-sectional estimates to their true values is an instance of nominal time. Nominal time is only applicable to the case of nonnegative data due to the importance of geometry in the mixed-signs case, so we confine its discussion to Part I. Part I covers nonnegative data. Part II relaxes the sign restriction to examine data of mixed sign. In each of these Parts, we specify the assumptions and create the simplest loss function that satisfies those assumptions and admits the property that the loss for a given relative difference (properly defined) increases in the initial value. A Lie symmetry permits elimination of an exponent. This enables each value of the remaining exponent to be associated with pairs of exponents lying on a line for the equivalent decompositions into power of the absolute difference and absolute relative difference (properly defined).
Part III concludes this article.
Part I Nonnegative Data
I.1 Introduction to Part I
Stock data, such as population, bank deposits and inventories are normally nonnegative data. Some flow data, such as births and deaths are also nonnegative. We develop loss function in Section I.2 to compare two sets of nonnegative values. We create four types of loss functions: unsigned, signed, time-invariant and a special case when the data are generated by processes sharing a particular set of criteria. Section I.3 discusses some applications, including general usage of loss functions, parametrizing loss functions from preexisting outlier criteria, using loss functions with GIS and parameterizing a variant of the loss functions of Section I.2 when the outliers are classified by another variable. These examples are based on actual Census Bureau applications.
I.2 The Loss Function
@@
We describe the assumptions used to generate the loss function $L(F;B)$ and its variants, where $F$ is the future value and $B$ is the base period value. The loss function is the penalty associated with the difference between $F$ and $B$. Roughly speaking, the greater the difference between $F$ and $B$, the greater the loss. Initially, $F$ is assumed to be one period after $B$. After we make the necessary assumptions in Subsection I.2.1, we specify the simplest forms of $L$ that also allow the values of the parameters of $L$ that make it increase in $B$ for a given relative difference. We then give the condition for the latter to happen. In Subsection I.2.2 we generalize $L$ to situations in which $F$ and $B$ may not be exactly one period apart. In Subsection I.2.3 we introduce the signed loss function for cases in which the sign of the difference is an additional important criterion. In Subsection I.2.4 we parametrize $L$ for comparing two sets of estimates of the same parameters. Throughout this paper, we assume $B$ and $F$ to be positive. Zeros, which frequently arise in the contexts in which we have worked, are either recoded to small values or omitted from the analysis.
I.2.1 The Unsigned Loss Function
@@
We construct the unsigned loss function $L$ by specifying three assumptions. [^5] The exposition in this Subsection through Property 1 (except for the use of Lie symmetry) follows Coleman (2025). The first assumption is that $L$ is symmetric in the differences:
Assumption I.1: $L(B+\epsilon,B)=L(B-\epsilon,B)$ for all $B,\epsilon\geq 0$.
@@
This assumption is not as innocuous as it looks. It is quite possible that, at least for some range of $B$, that positive and negative differences have differential impacts. However, the resulting asymmetry complicates the definition of $L$. We relax this assumption in Subsection I.2.3 by developing the signed loss function, which allows the possibility of incorporating the direction of the difference $\epsilon$. The symmetry of $L$ allows us to use the equivalent notation $\ell(\epsilon,B)=L(F;B)$ where $\epsilon=|F-B|$.
@@
Assumption I.2: $\partial\ell/\partial\epsilon>0$ for all $\epsilon\geq 0$.
@@
Finally, we want $L$, or, equivalently, $\ell$, to decrease in $B$. This means that for a given value of $\epsilon$, the loss associated with it decreases with its associated initial value. This has two justifications. First, for example, a difference of 500 when the initial value is 1,000 is a whopping 50%, a highly significant difference. However, the same difference, when the initial value is 1,000,000 is akin to a roundoff error. Second, when performing estimates or taking samples, the coefficient of variation $\sigma/\mu$, where $\sigma$ is the standard deviation and $\mu$ is the expected value, decreases in $B$. Coleman’s experience is that all areas tend to have about the same roundoff errors. Again, these are proportionately greater in small areas. We state this formally as:
Assumption I.3: $\partial\ell/\partial B<0$, or, equivalently, $\partial L/\partial B<0$ for all $B>0$.
The simplest function which satisfies Assumptions I.1–I.3 and admits Property I.1 below is the Cobb-Douglas function
$\epsilon^{p}B^{q}$ (I.1a)
@@
$|F-B|^{p}B^{q}.$ (I.1b)
An observed pair $(F;B)$ is an outlier whenever $L(F;B)>C$, where $C$ is a predetermined critical value. [^6] CC can also be determined from the data by taking a predetermined quantile or a multiple of the interquartile range of $L$ (Tukey 1977). We refer to outliers as being critical. We also refer to the equation $L(F;B)=C$ as the equation of criticality. The choice of $p,q$ and $C$ is an empirical matter. Only a practitioner’s experience with data can determine when data are suspect and incorporate these suspicions into parameters. One thing to note is that the loss function is ordinal: raising $L$ or $\ell$ and $C$ to any positive power $m$ leaves the rankings of losses unchanged. It is only the rankings of losses that are important. Formally, the transformation $p\to p_{m}=mp$, $q\to q_{m}=mq$ and $C\to C_{m}=mC$ creates a new equation of the same form. The same independent variables create the same criticality (or lack thereof). This is called a Lie symmetry or Lie invariance. The upshot is that one parameter can be fixed and the other two varied to generate new loss functions and critical values. In subsequent development, we set $p$ equal to 1. Thus,
$\ell(F,B)=\epsilon B^{q}$ (I.2a)
@@
Property I.1: The loss function defined by equations (I.2a) and (I.2b) increases in $B$ for any given absolute relative difference. This is assured whenever q>1q>-1.
The reader may notice that $q=0$ turns the first terms of equations (I.2a) and (I.2b) into the absolute values of the differences. Thus, values of $q$ between 0 and 1–1 represent various tradeoffs of absolute differences and absolute relative differences. Consider the product of the $r$th power of the absolute difference and the $s$th power of the absolute relative difference, where $r,s>0$,
$|F-B|^{r}\left(\frac{|F-B|}{B}\right)^{s}.$ (I.4)
By the Lie symmetry invoked above, loss function (I.4) is isomorphic to the loss function $|F-B|B^{-s/r+s}$. Thus, any value of $q$ corresponds to an infinite number of pairs $(r,s)$ where q=s/(r+s)q=–s/(r+s). Geometrically, the same loss function is generated for all $(r,s)$ lying on the line r=(1+q)sr=–(1+q)s.
Subsection I.3.4 provides an example of a variant of this loss function approach when the outliers are classified by another variable. Changes in the variable of interest are declared outliers contingent on another variable. Assumption I.3 is violated, but, in this example, Property I.1 holds.
I.2.2 The Time-Invariant Loss Function
@@
$L(F_{it};B_{i},t)=\left(\frac{|F_{it}-B_{i}|}{B_{i}}\right)^{\frac{1}{t}}B_{i}^{q+1}=|F_{it}-B_{i}|^{\frac{1}{t}}B_{i}^{q+1+\frac{1}{t}}$ (I.5)
Equation (I.5) puts the geometric average absolute relative difference on the same basis for all $t$.
Now, we can use Lie symmetry to make the exponent of $|F_{it}-B_{i}|$ unity by multiplying both exponents by $t$.
$L(F_{it};B_{i},t)=|F_{it}-B_{i}|B_{i}^{tq+t-1}.$ (I.6)
@@
I.2.3 The Signed Loss Function
@@
The time-invariant signed loss function is
$S(F_{it};B_{i},t)=(F_{it}-B_{i})B_{i}^{tq+t-1}.$ (I.8)
I.2.4 Comparing Two Sets of Data: A Specialization of the Loss Function
Suppose that the sets $\mathbf{B}=\{B_{i}\}$ and $\mathbf{F}=\{F_{i}\}$ represent two versions of estimates of the true values $\mathbf{A}=\{A_{i}\}$. [^7] The derivations in this Subsection were done with assistance from ChatGPT 5. Suppose that both the $B_{i}$ and $F_{i}$ are unbiased estimators of the $A_{i}$ and that their variances are proportionate to the $A_{i}$ (i.e., $\operatorname{Var}(B_{i})=\operatorname{Var}(F_{i})=\sigma^{2}A_{i}$.) [^8] This can be justified by assuming that the $B_{i}$ and $F_{i}$ are both sums of $A_{i}$ independent (or, more weakly, jointly uncorrelated) estimators with expected value 1 and variance $\sigma^{2}$. Suppressing subscripts, consider the expected value of the squared difference of $B$ and $F$, $\operatorname{\mathbf{E}}(F–B)^{2}$:
$\operatorname{\mathbf{E}}(F-B)^{2}=\operatorname{Var}(F)+\operatorname{Var}(B)-2\operatorname{Cov}(F,B)+(\operatorname{\mathbf{E}}F-\operatorname{\mathbf{E}}B)^{2}$ (I.9)
@@ I.2.4 Comparing Two Sets of Data: A Specialization of the Loss Function
$\operatorname{\mathbf{E}}(F-B)^{2}=\operatorname{Var}(F)+\operatorname{Var}(B)-2\operatorname{Cov}(F,B).$ (I.10)
@@
$\operatorname{\mathbf{E}}(F-B)^{2}=2\sigma^{2}A-2\operatorname{Cov}(F,B).$ (I.11)
@@
$\operatorname{Cov}(F,B)=\rho\operatorname{Var}(F)\operatorname{Var}(B)=\rho\sigma^{2}A.$ (I.12)
Substitutiong equation (I.12) into equation (I.11) yields
$\operatorname{\mathbf{E}}(F-B)^{2}=2(1-\rho)\sigma^{2}A\propto A.$ (I.13)
@@
$(F-B)^{2}\propto A=\operatorname{\mathbf{E}}B.$ (I.14)
@@
$|F-B|B^{-\frac{1}{2}}\propto 1.$ (I.16)
This means that, in this situation, we can use the loss functions (I.2a) and (I.2b) with q=1/2q=–1/2. Since the null distributions of $B$ and $F$ are assumed unknown, it is impossible to do any significance testing. Moreover, since we are dealing with the entire population, sampling theory breaks down. Even though estimates of $\rho$ and $\sigma$ can be computed from the data, it is unclear how to use normal (or any other distributional) theory for testing. The simplest thing to do in this situation is to rank the losses in descending order and empirically decide which critical value to use.
Again, the signed loss function (I.7) can be used with q=1/2q=–1/2.
I.3 Applications
We illustrate the use of loss functions by first outlining a general procedure for using loss functions in Subsection I.3.1. Next, three different examples of loss functions are shown. In the first example, in Subsection I.3.2, preexisting outlier criteria in terms of critical ratios by size class are transformed into a loss function. The second example, in Subsection I.3.3, uses real-world data and GIS to compare the 2010 and 2020 Census county counts using the q=1/2q=–1/2 loss function of Subsection I.2.4 as well as absolute and absolute relative differences to evaluate differences between these two sets of estimates. The third example, in Subsection I.3.4, uses preexisting outlier criteria for a variable, which are classified by another variable, to generate a different type of loss function.
I.3.1 General Procedure for Using Loss Functions
Loss function evaluations usually begin by recoding zeros to a small positive value (the exact value determined by the range and smallest value of the data) and setting q=.5q=-.5. If time is chronological, the analyst then has to examine the data and the rankings of their associated losses. If, in their opinion, too many observations with small changes occurring to small base values are ranked highly, then $q$ should be increased. If, on the other hand, too many observations with small changes to large base values are ranked highly, then $q$ should be decreased. This process continues until the analyst is satisfied with the loss rankings. Our experience has found that initially changing $q$ by .1 is satisfactory. If time is nominal, our experience is that two comparisons should be done using both variables for $B$.